# Context-Selective State Space Models: Feedback is All You Need

### v1

@@ Abstract
Transformers, powered by the attention mechanism, are the backbone of most foundation models, yet they suffer from quadratic complexity and difficulties in dealing with long-range dependencies in the input sequence. Recent work has shown that state space models (SSMs) provide an efficient alternative, with the S6 module at the core of the Mamba architecture achieving state-of-the-art results on long-sequence benchmarks. In this paper, we introduce the COFFEE (COntext From FEEdback) model, a novel time-varying SSM that incorporates state feedback to enable context-dependent selectivity, while still allowing for parallel implementation. Whereas the selectivity mechanism of S6 only depends on the current input, COFFEE computes it from the internal state, which serves as a compact representation of the sequence history. This shift allows the model to regulate its dynamics based on accumulated context, improving its ability to capture long-range dependencies. In addition to state feedback, we employ an efficient model parametrization that removes redundancies present in S6 and leads to a more compact and trainable formulation. On the induction head task, COFFEE achieves near-perfect accuracy with two orders of magnitude fewer parameters and training sequences compared to S6. On MNIST, COFFEE largely outperforms S6 within the same architecture, reaching 97% accuracy with only 3585 parameters. These results showcase the role of state feedback as a key mechanism for building scalable and efficient sequence models.
@@ 1 Introduction
Recurrent Neural Networks and Long Short-Term Memory Networks have been the dominant architectures for sequence modeling up to the introduction of Transformers [1]. The latter are the backbone of modern foundation models, namely large-scale models that are pre-trained on massive amounts of data (text, audio, video, etc.) and subsequently fine-tuned, using task-specific data, for specific downstream tasks. By leveraging the celebrated attention mechanism, transformers achieved a groundbreaking improvement in generative Artificial Intelligence (AI), [2, 3, 4, 5, 6]. The attention mechanism, though revolutionary, suffers from some limitations. In particular, a quadratic time complexity with respect to the length of the input sequence and severe difficulties in capturing long-range dependencies call for further research and improvements. To this end, many variations of the attention mechanism have been proposed; see [7, 8, 9, 10, 11]. A completely different and extremely promising approach is the use of linear State-Space Models (SSMs): introduced in the 1960’s, they quickly revolutionized control and automation engineering [12, 13], and their use spread to neighboring fields, from biology [14] to economics [15] and statistical modeling [16]. In fact, a significant stream of valuable contributions have recently appeared that replace the attention mechanism with SSMs [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Other interesting works, such as [30] and [31] employ nonlinear SSMs to address questions about asymptotic behavior of Transformers as the number of encoders layers increases.
@@
The state of the art of SSM in sequence modeling is called S6 and it lies at the core of the so-called Mamba architecture [20]. The latter exhibits great capabilities and is able to outperform the state-of-the-art transformer in challenging tasks, including the Long Range Arena (LRA), see [20] and [32]. The reason for this success lies in the fact that the state is, by definition, the mathematical object that encodes all the relevant information on the past evolution of a system needed to predict its future behavior [^1] See Section 2 and Appendix A for more details on this.. This separation property allows states to store elusive long-range dependencies. Of course, the main challenge is to derive a universal class of state-space models whose parameters can be learned for each specific task. The S6 module captures, or rather approximates, this universality by resorting to (linear) time-varying models.
@@
In this paper, we propose a further dramatic improvement: a novel class of SSMs that we call COFFEE (COntext From FEEdback) which is endowed with feedback capabilities. Dynamical systems endowed with feedback loops are at the basis for the large majority of complex natural and artificial phenomena, as they provide a system with the possibility of changing its behavior on the basis of its current state, which, as previously recalled, contains all the relevant information about the past of the sequence under analysis. In this way, we are able to induce context-selectivity in the behavior of COFFEE: a property of paramount importance, which appears to give a momentous boost to the system’s performance. With respect to S6, COFFEE has another advantage (which could also be implemented in the S6 model): by exploiting a change of basis in the state-space representation of the model, we are able to reduce appreciably the number of learnable parameters, offering an obvious edge in terms of both numerical implementation and mechanistic interpretability.
@@
We tested COFFEE in several versions of the induction head benchmark and in the MNIST benchmark. In these tasks, our model appears to perform significantly better than the state-of-the-art S6 in more than one way. In fact, not only does COFFEE give much better results in terms of the accuracy of its output, but it also requires a smaller number of learnable parameters and a much smaller training set to converge. The closed-loop COFFEE model is obtained by a nonlinear state-feedback action on a time-varying linear model. In general, this kind of structure leads to computational drawbacks because it is not amenable to parallelization and hence cannot take advantage of the highly parallel architecture of modern GPU’s. In COFFEE, however, the feedback structure is constrained in such a way that, on the one hand, the performance degradation (with respect to the case of a more general feedback structure) is negligible and, on the other, the Jacobian of the state transition function is guaranteed to be diagonal. As a consequence, we can leverage recent results in [33] to obtain scalable parallel training that is computationally highly efficient.
@@
The remainder of this paper is organized as follows. In Section 2 we give a brief background on state-space models, with further details and useful proofs provided in A. Based on the review of state space models, in Section 3, we present our model along with its mathematical motivations. In Section 4, we prove the effectiveness of our model by comparing its performance with the state-of-the-art S6 model [20], in both the induction head and MNIST tasks. In Section 5, we provide some theoretical insights, based on an analysis made on a simplified version of the induction head task, whose details are presented in Appendix D. We summarize our findings in Section 6, where we also outline the next steps in the development and testing of the model.
2 Background on SSMs
@@ 2.1 Linear, discrete SSM
State space models are powerful tools to model, analyze and control dynamical systems [13]. These are models in which the relation between the input or control sequence $u(k)\in\mathbb{R}^{m}$ and the sequence $y(k)\in\mathbb{R}^{p}$ associated to the quantities of interest for the system at hand and called the output, is mediated by a signal $x(k)\in\mathbb{R}^{n}$, which is called the state of the system. For more details, see Appendix A. In this paper, we leverage the class of linear, time-varying, discrete-time systems with finite-dimensional state space. The models of this class are described by a couple of vector equations of the form
@@
where the matrix $A(k)\in\mathbb{R}^{n\times n}$ is called state transition matrix, $B(k)\in\mathbb{R}^{n\times m}$ is called input matrix, $C(k)\in\mathbb{R}^{p\times n}$ is called output matrix. From (1) it follows that, given the state $x(k)$ at the present “time” $k$, the past of the input is irrelevant for the future of the output sequence. This key feature of the state is known as the separation property, and is the fundamental reason why the state functions as memory. The state can then be considered as a natural representative of the context set by the past data, and as such will be exploited in our setting in selecting the dynamics conveniently.
@@ 2.2 Relevance of SSMs to Sequence Modeling
In machine learning, especially in sequence modeling tasks (e.g. time series forecasting, language modeling), SSMs offer several advantages:
Efficiency: The convolutional formulation allows for fast evaluation via the Fast Fourier Transform (FFT), reducing the complexity from $\mathcal{O}(L^{2})$ (as in the attention mechanism) to $\mathcal{O}(L\log L)$, $L$ being the length of the input signal $u(k)$;
Long-range Dependencies: By construction, the state of the system contains a compressed representation of the system’s past history. In the context of sequence modeling, the input $u(k)$ corresponds to the tokens provided to the model, meaning that the state is capable of “remembering” all previously seen tokens in a compact form;
@@
Recent neural architectures such as S4 [19] and Mamba [20] leverage these ideas, combining them with modern training techniques and nonlinearities to build powerful sequence models. In particular, S4 uses a time-invariant model, namely a model of the form (1) where, however, the matrices $A,B,C$ are independent of $k$, while the main novelty of S6, the core module of Mamba, is that it uses time-varying matrices $A(k),B(k),C(k)$.
In the context of sequence modeling, a state-space model as in Equation (1) is used as black box in foundation models. The exogenous input $u(k)$ represents the $k$-th input token fed into the model, the state $x(k)$ stores relevant information about the tokens up the $k$-th one, while $y(k)$ is the $k$-th output: for generative or prediction models $y(k)$ is used for the prediction of $u(k+1)$; more generally, $y(k)$ is the signal used to solve the task at hand.
In general, the matrices $A,B,C$ are learned using a stochastic gradient descent. Computational efficiency and initialization are crucial in this context, therefore, the matrices $A,B,C$ often have a specific structure, giving rise to the concept of structured state space models. For example, the matrix $A$ is typically chosen to be diagonal.
@@ 3 COFFEE: Towards Context-Selective SSMs
In this section, we introduce a novel state space model which we call COFFEE (COntext From FEEdback). The state-of-the-art S6 SSM, proposed in [20], enables selectivity conditioned on the current token by adapting the system dynamics to the current input. In contrast, the main novelty of COFFEE is to replace this token-based selectivity with context-based selectivity, achieved through a feedback mechanism that modulates the dynamics based on the state of the model.
Both COFFEE and S6 are implemented within the core module architecture described in Appendix E, which maps sequences over a finite vocabulary $V$ into embeddings in $\mathbb{R}^{D}$, and processes each embedding feature with a dedicated single-input single-output (SISO) state space model. To motivate our modifications, we now briefly review the structure of the S6 SISO component (a more detailed analysis is available in Appendix B).
@@ 3.1 Background: The S6 Model
As mentioned above, in S6 each feature $u(k)_{i}$ of the embedding $u(k)\in\mathbb{R}^{D}$ is handled independently by a dedicated SISO state-space model $\Sigma^{(i)}$. This design allows the architecture to scale across the embedding dimension by running $D$ parallel SISO models.
The discrete-time update for each feature is:
$\begin{cases}x(k)^{(i)}=e^{A^{(i)}\Delta_{i}(k)}x(k-1)^{(i)}+\big(A^{(i)}\big)^{-1}\!\big(e^{A^{(i)}\Delta_{i}(k)}-I\big)B(k)\,u(k)_{i},\\ y(k)^{(i)}=C(k)\,x(k)^{(i)},\end{cases}$ (2)
where $x(k)^{(i)}\in\mathbb{R}^{n}$ is the hidden state of the $i$-th feature, $A^{(i)}=\mathrm{diag}(\lambda_{0}^{(i)},\dots,\lambda_{n-1}^{(i)})\in\mathbb{R}^{n\times n}$ is a diagonal transition matrix, and $u(k)_{i}$ denotes the $i$-th feature of the embedding vector $u(k)$.
The input-to-state and state-to-output mappings are defined as
$B(k)=W_{B}u(k),\qquad C(k)=(W_{C}u(k))^{\top},$
with $W_{B},W_{C}\in\mathbb{R}^{n\times D}$ shared across all features. Similarly, the step sizes $\Delta_{i}(k)$ are determined from the current input as the components of the vector
@@
where $W_{D}\in\mathbb{R}^{D\times D}$ and $\zeta(\cdot)$ denotes the softplus function.
A distinctive aspect of S6 is its token-based selectivity: the parameters $B(k)$, $C(k)$, and $\Delta(k)$ depend directly on the current token $u(k)$. This enables the model to dynamically amplify or suppress contributions from different inputs, effectively filtering out irrelevant or noisy symbols while focusing on semantically meaningful ones. Such input-conditioned adaptation is particularly advantageous in sequential domains (e.g., speech, where disfluencies such as “uh” carry little meaning). This token-dependent, time-varying structure is widely regarded as the core innovation that underpins the strong empirical performance of Mamba and its descendants.
@@ 3.2 The COFFEE Model
The central novelty of COFFEE is the use of state feedback to achieve context-based selectivity. Unlike S6, where the update gates $\Delta_{i}(k)$ depend only on the current input token, COFFEE computes them as functions of the internal state, allowing the model to adapt its dynamics based on the accumulated context. It is well known and very intuitive that open-loop systems are fragile and sensitive to model uncertainties and disturbances, whereas closed-loop feedback ensures robustness; COFFEE exploits this principle to achieve context-based selectivity, improving robustness and coherence in generative modeling.
@@
In S6, the step size $\Delta_{i}(k)$ controls the tradeoff between retaining past state and incorporating new input. However, its action is embedded within matrix exponentials, which complicates direct analysis. To make this dependence more explicit and obtain a more interpretable formulation, we approximate $e^{A^{(i)}\Delta_{i}(k)}$ with its first-order Taylor expansion around $\Delta_{i}(k)=0$. This yields the dynamics:
$\begin{cases}x(k)^{(i)}=\big(I+A^{(i)}\Delta_{i}(k)\big)\,x(k-1)^{(i)}+\Delta_{i}(k)B^{(i)}u(k)_{i},\\ y(k)^{(i)}=C^{(i)}x(k)^{(i)},\end{cases}$ (4)
where $B^{(i)}\in\mathbb{R}^{n\times 1}$ and $C^{(i)}\in\mathbb{R}^{1\times n}$ are time-invariant, feature-specific parameters. The latter choice on the input and output matrices is motivated by the fact that, due to discretization, the input-to-state mapping in (2) is already feature-specific, while we additionally assign feature-specific $C^{(i)}$ to maintain symmetry and let selectivity be governed solely by the step size $\Delta_{i}(k)$.
By replacing exponentials with affine dependence on $\Delta_{i}(k)$, the simplified model in (4) preserves the essential gating mechanism of S6 while rendering the dynamics simpler and more amenable to mechanistic interpretation.
@@
A central strength of S6 lies in its token-based selectivity: the update gates $\Delta_{i}(k)$ are computed as functions of the current input token $u(k)$, so that the model can suppress irrelevant inputs and amplify informative ones. However, token-based selectivity is limited because identical tokens may play very different roles depending on context. For example, if we ask for an approximation of the number “111.11”, where each digit “1” is represented by the same token, the model must treat different occurrences of “1” differently depending on their position.
To address this limitation, we observe that the hidden state $x(k-1)^{(i)}$ of each subsystem $\Sigma^{(i)}$ already encodes a compressed representation of the full history up to step k1k-1, and therefore provides a natural source of contextual information. We therefore replace the token-based gating (3) with a state-dependent gating rule:
@@
where $w_{D}^{(i)}\in\mathbb{R}^{n}$ is a learnable parameter vector, $\odot$ denotes the Hadamard product, and $\sigma(\cdot)$ is the sigmoid function applied componentwise to vectors. [^2] We use the sigmoid rather than the softplus, since its bounded range keeps $\Delta_{i}(k)$ limited, helping maintain the validity of the Taylor approximation in (4). By doing so, each subsystem adapts its update gate from its past history rather than the raw input token, enabling context-based selectivity: identical tokens can be filtered differently depending on their role in the evolving sequence. Concretely, the gating vector $\Delta_{i}(k)\in\mathbb{R}^{n}$ enters the dynamics in (4) as a diagonal state feedback matrix, which modulates the contribution of the past state through
$(I+A^{(i)}\mathrm{diag}(\Delta_{i}(k)))x(k-1)^{(i)},$
and the input through
$\mathrm{diag}(\Delta_{i}(k))B^{(i)}u_{i}(k).$
The resulting subsystem dynamics is therefore context-adaptive, taking the form
@@
As a final step, we note that the parameterization of $(B^{(i)},C^{(i)})$ in (6) exhibits parameter redundancy: through suitable changes of basis (corresponding to simple rescalings) the same input-output behavior can be preserved while reducing the number of parameters. This fact is formalized in the following proposition, whose proof is deferred to the Appendix (see Section C.3 where a few comments are also provided on the proposition assumptions).
@@
Assume that at least one embedding vector has all the entries that are non-zero and that all the entries of the $B^{(i)}$ are non-zero. For the family of subsystems $\{\Sigma^{(i)}\}_{i=0}^{D-1}$ defined in Eq. (6), there exist changes of basis in the input and state spaces such that the structure of Eq. (6) is preserved and:
  1. (i)
    all input-to-state matrices $B^{(i)}$ can be fixed to the constant vector $\mathbf{1}=[1,\dots,1]^{\top}$, with their scaling absorbed into the corresponding $C^{(i)}$, and
  2. (ii)
    one embedding vector can be normalized to $\mathbf{1}$.
The transformations in Proposition 1 preserve both the input-output map and the diagonal structure of $A^{(i)}$. Consequently, the input-to-state matrices $B^{(i)}$ are redundant, as their effect can always be absorbed into $C^{(i)}$, and one embedding vector can be fixed to $\mathbf{1}$. This reduces the parameter count by $nD+D$ without loss of expressivity, yielding a leaner architecture that is easier to train.

### v2

@@ Abstract
Transformers, powered by the attention mechanism, are the backbone of most foundation models, yet they suffer from quadratic complexity and difficulties in dealing with long-range dependencies in the input sequence. Recent work has shown that state space models (SSMs) provide a promising alternative. In this paper, we introduce the COFFEE (COntext From FEEdback) model, a novel time-varying SSM that incorporates state feedback to enable context-dependent selectivity, while still allowing for parallel implementation. This idea allows the model to regulate its dynamics based on the context described by the internal state, which embodies a compact representation of the input history. State feedback allows COFFEE to improve its ability to capture long-range dependencies: on the induction head task, it achieves near-perfect accuracy with two orders of magnitude fewer parameters and training sequences compared to S6 (the SSM of Mamba). On MNIST, COFFEE largely outperforms S6 within the same architecture, reaching 97% accuracy with only 3585 parameters. These results showcase the role of state feedback as a key mechanism for building scalable and efficient sequence models.
@@ 1 Introduction
The transformer based on the celebrated attention mechanism, [42], is the key of a recent groundbreaking improvement in generative Artificial Intelligence (AI), [9, 1, 17, 41, 12]. Yet, the attention mechanism suffers from some limitations due to a quadratic time complexity with respect to the length of the input sequence and severe difficulties in capturing long-range dependencies. Several variations have been proposed to mitigate this problems, [4, 44, 25, 6, 21]. A completely different and extremely promising approach is the use of linear State-Space Models (SSMs): introduced in the 1960’s, they quickly revolutionized control and automation engineering [23] and many neighboring fields. In fact, a significant stream of valuable contributions have recently appeared that replace the attention mechanism with SSMs [39, 18, 16, 15, 34, 19, 7, 22, 31, 35, 36, 32]. In [45] a unifying framework is presented which embeds the transformer and many SSMs into the general setting of stochastic realization, [28]; In this framework a more general model is also presented which couples the advantages of the transformer short-term eidetic memory with those of the long-term fading memory of SSMs. The connections between LLMs and control systems has been thoroughly investigated. In particular, controllability is analyzed in [40], observability in [29] and stability in [30].
@@
The starting point of this paper is the so called S6 SSM which lies at the core of the so-called Mamba architecture [15]. [^1] The evolutions Mamba-2 [8], and Mamba-3 are powered by similar SSMs. The latter exhibits great capabilities and is able to outperform the state-of-the-art transformer in challenging tasks, including the Long Range Arena (LRA), see [15] and [2]. The reason for this success lies in the fact that the state is, by definition, the mathematical object that encodes all the relevant information on the past evolution of a system needed to predict its future behavior [^2] See Section 2 and Appendix A for more details on this.. Of course, the main challenge is to derive a universal class of state-space models whose parameters can be learned for each specific task. The S6 module captures, or rather approximates, this universality by resorting to (linear) time-varying models.
@@
In this paper, we propose a novel class of SSMs that we call COFFEE (COntext From FEEdback) which is endowed with feedback capabilities. Dynamical systems endowed with feedback loops are at the basis for the large majority of complex natural and artificial phenomena, as they enable a system to adapt its behavior based on its current state, which, as previously recalled, contains all the relevant information about the past of the sequence under analysis. In this way, we are able to induce context-selectivity in the behavior of COFFEE: a property of paramount importance, which appears to yield substantial performance improvements in model performance. With respect to S6, COFFEE has another advantage (which could also be implemented in the S6 model): by exploiting a change of basis in the state-space representation of the model, we are able to reduce appreciably the number of learnable parameters, offering an obvious edge in terms of both numerical implementation and mechanistic interpretability.
@@
We tested COFFEE in several versions of the induction head task and in the MNIST benchmark. Across these tasks, our model consistently outperforms S6 along multiple dimensions. In particular, COFFEE not only achieves substantially higher output accuracy, but also requires fewer learnable parameters and significantly smaller training datasets to reach convergence. The closed-loop COFFEE model is obtained by a nonlinear state-feedback action on a time-varying linear model. In general, this kind of structure leads to computational drawbacks because it is not amenable to parallelization and hence cannot take advantage of the highly parallel architecture of modern GPUs. In COFFEE, however, the feedback structure is constrained in such a way that, on the one hand, the performance degradation (with respect to the case of a more general feedback structure) is negligible and, on the other, the Jacobian of the state transition function is guaranteed to be diagonal. As a consequence, we can leverage recent results in [13] to obtain scalable and highly efficient parallel training.
Finally, we mention that [10] proposes a biologically-inspired nonlinear SSM (LrcSSM) with diagonal Jacobian. While LrcSSM relies on intrinsically nonlinear state- and input-dependent dynamics, COFFEE restricts nonlinearity to a diagonal state-feedback gate that modulates a fixed linear SSM backbone, resulting in a more constrained and parameter-efficient model.
@@
The remainder of this paper is organized as follows. In Sec. 2 we give a brief background on SSMs, with further details provided in Appendix A. In Sec. 3, we present our model along with its mathematical motivations. In Sec. 5, we assess the effectiveness of COFFEE by comparing it with S6 [15] on the induction head and MNIST tasks. In Sec. 6, we provide some theoretical insights, based on an analysis made on a simplified version of the induction head task, whose details are presented in Appendix E. We summarize our findings in Sec. 7, where we also outline the next steps in the development and testing of the model.
2 SSMs and Sequence Modeling
@@ 2.1 Linear, discrete SSM
State space models are powerful tools to model, analyze and control dynamical systems [23]. These are models in which the relation between the input or control sequence $u(k)\in\mathbb{R}^{m}$ and the sequence $y(k)\in\mathbb{R}^{p}$ associated to the quantities of interest for the system at hand and called the output, is mediated by a signal $x(k)\in\mathbb{R}^{n}$, which is called the state of the system. For more details, see Appendix A. In this paper, we leverage the class of linear, time-varying, discrete-time systems with finite-dimensional state space. The models of this class are described by a pair of vector equations of the form
@@
where the matrix $A(k)\in\mathbb{R}^{n\times n}$ is called state transition matrix, $B(k)\in\mathbb{R}^{n\times m}$ is called input matrix, $C(k)\in\mathbb{R}^{p\times n}$ is called output matrix. From (1) it follows that, given the state $x(k)$ at the present “time” $k$, the past of the input is irrelevant for the future of the output sequence. This key feature of the state is known as the separation property, and is the fundamental reason why the state functions as memory. The state can then be considered as a natural representative of the context set by the past data, and will be exploited in our setting in selecting the dynamics conveniently.
@@ 2.2 Relevance of SSMs to Sequence Modeling
In the context of sequence modeling, a state-space model as in (1) is used as a black-box in foundation models. The exogenous input $u(k)$ represents the $k$-th input token fed into the model, the state $x(k)$ stores relevant information about the tokens up to the $k$-th one, while $y(k)$ is the $k$-th output: for generative or prediction models $y(k)$ is used to predict $u(k+1)$; more generally, $y(k)$ is the signal used to solve the task at hand. This offers several advantages:
Efficiency: SSMs are based on linear recurrences that enable efficient inference and allow training to be parallelized via FFT-based or associative scan algorithms.
Long-range dependencies: By construction, the state of the system contains a compressed representation of the past history of the system: the state is capable of “remembering” all previously seen tokens in a compact form;
@@
Recent neural architectures such as S4 [16] and S6 [15] leverage these ideas, combining them with modern training techniques and nonlinearities to build powerful sequence models. In particular, S4 uses a time-invariant model, namely a model of the form (1) where, however, the matrices $A,B,C$ are independent of $k$, while the main novelty of S6 is that the matrices of the SSM are input-dependent and hence time-varying.
In general, the matrices $A,B,C$ are structured (typically $A$ is diagonal) and are learned during training.
@@ 3 COFFEE: Towards Context-Selective SSMs
In this section, we introduce a novel state space model which we call COFFEE (COntext From FEEdback). S6 [15] enables selectivity conditioned on the current token by adapting the system dynamics to the current input. COFFEE replaces this token-based selectivity with context-based selectivity, achieved through a feedback mechanism that modulates the dynamics based on the state of the model.
Both COFFEE and S6 are implemented within the core module architecture described in Appendix F, which maps sequences over a finite vocabulary $V$ into embeddings in $\mathbb{R}^{D}$, and processes each embedding feature with a dedicated single-input single-output (SISO) state-space model. To motivate our modifications, we now briefly review the structure of the S6 SISO component (a more detailed analysis is available in Appendix B).
@@ 3.1 Background: The S6 Model
As mentioned above, in S6 each feature $u(k)_{i}$ of the embedding $u(k)\in\mathbb{R}^{D}$ is handled independently by a dedicated SISO state-space model $\Sigma^{(i)}$. This design allows the architecture to scale across the embedding dimension by running $D$ parallel SISO models. The SSM for each feature is:
$\begin{cases}x(k)^{(i)}=e^{A^{(i)}\Delta_{i}(k)}x(k-1)^{(i)}+\ \big(A^{(i)}\big)^{-1}\!\big(e^{A^{(i)}\Delta_{i}(k)}-I\big)B(k)\,u(k)_{i}\\ y(k)^{(i)}=C(k)\,x(k)^{(i)}\end{cases}$ (2)
where $x(k)^{(i)}\in\mathbb{R}^{n}$ is the hidden state of the $i$-th feature, $A^{(i)}=\mathrm{diag}(\lambda_{0}^{(i)},\dots,\lambda_{n-1}^{(i)})\in\mathbb{R}^{n\times n}$ is a diagonal matrix (a suitable learning implementation is adopted to avoid vanishing $\lambda_{j}^{(i)}$), and $u(k)_{i}$ denotes the $i$-th feature of the embedding vector $u(k)$. The input-to-state and state-to-output mappings are $B(k)=W_{B}u(k)$ and $C(k)=(W_{C}u(k))^{\top}$, with $W_{B},W_{C}\in\mathbb{R}^{n\times D}$ shared across all features. The step sizes $\Delta_{i}(k)$ are determined as the components of the vector
@@
where $W_{D}\in\mathbb{R}^{D\times D}$ and $\zeta(\cdot)$ denotes the softplus function. Notice that, while $B(k)$ is the same for each $i$, the overall input matrix $(A^{(i)})^{-1}\!(e^{A^{(i)}\Delta_{i}(k)}-I)B(k)$ is indeed feature-dependent. A distinctive aspect of S6 is its token-based selectivity: the parameters $B(k)$, $C(k)$, and $\Delta(k)$ depend directly on the current token $u(k)$. This enables the model to dynamically amplify or suppress contributions from different inputs, effectively filtering out irrelevant symbols while focusing on meaningful ones. Such input-conditioned adaptation is particularly advantageous in sequential domains (e.g., speech, where disfluencies carry little meaning). This token-dependent, time-varying structure is widely regarded as the core innovation that underpins the strong empirical performance of Mamba and its descendants.
@@ 3.2 The COFFEE Model
The central novelty of COFFEE is the use of state feedback to achieve context-based selectivity. Unlike S6, where the update gates $\Delta_{i}(k)$ depend only on the current input token, COFFEE computes them as functions of the internal state, allowing the model to adapt its dynamics based on the accumulated context. It is well known that open-loop systems are fragile and sensitive to model uncertainties and disturbances, whereas closed-loop feedback ensures robustness; COFFEE exploits this principle to achieve context-based selectivity, leading to more stable and consistent generative behavior.
@@
In S6, the step size $\Delta_{i}(k)$ controls the tradeoff between retaining past state and incorporating new input. However, its action is embedded within matrix exponentials, which complicates direct analysis. To make this dependence more explicit and obtain a more interpretable formulation, we approximate $e^{A^{(i)}\Delta_{i}(k)}$ with its first-order Taylor expansion around $\Delta_{i}(k)=0$. Furthermore, we replace the matrices $B(k)$ and $C(k)$ with feature-dependent, time-invariant ones. The choice of feature-dependent matrices is motivated by the fact that, due to discretization, the input-to-state mapping in (2) is already feature-dependent, while we additionally assign feature-dependent $C^{(i)}$ to maintain symmetry. Instead, time-invariance allows selectivity be governed solely by the step size $\Delta_{i}(k)$. This yields the dynamics:
$\begin{cases}x(k)^{(i)}=\big(I+A^{(i)}\Delta_{i}(k)\big)\,x(k-1)^{(i)}+\ \Delta_{i}(k)B^{(i)}u(k)_{i}\\ y(k)^{(i)}=C^{(i)}x(k)^{(i)}\end{cases}$ (4)
where $B^{(i)}\in\mathbb{R}^{n\times 1}$ and $C^{(i)}\in\mathbb{R}^{1\times n}$ are time-invariant, feature-dependent matrices. By replacing exponentials with affine dependence on $\Delta_{i}(k)$, the simplified model in (4) preserves the essential gating mechanism of S6 while rendering the dynamics simpler and more amenable to interpretation.
@@
A key strength of S6 lies in its token-based selectivity: the update gates $\Delta_{i}(k)$ are computed as functions of the current input token $u(k)$, so that the model can suppress irrelevant inputs and amplify informative ones. However, token-based selectivity is inherently limited as identical tokens may play different roles depending on context. As a simple example, consider approximating the number “111.11”, where each digit “1” is represented by the same token. In this case, the model must treat different occurrences of “1” differently depending on their position. To address this limitation, we observe that the hidden state $x(k-1)^{(i)}$ of each subsystem $\Sigma^{(i)}$ already encodes a compressed representation of the full history up to step k1k-1, and thus provides a natural source of contextual information. We therefore replace the input-dependent gating mechanism with a state-dependent one. In addition, we extend the scalar gate $\Delta_{i}(k)$ in (4) to a vector-valued gate to enable efficient parallel training, as discussed in Sec. 4. Concretely, we replace the scalar token-based gating (3) with a vector-valued, state-dependent gating rule:
@@
where $w_{D}^{(i)}\in\mathbb{R}^{n}$ is a learnable parameter vector, $\odot$ denotes the Hadamard product, and $\sigma(\cdot)$ is the sigmoid function applied componentwise to vectors. [^3] We use the sigmoid rather than the softplus, since its bounded range keeps $\Delta_{i}(k)$ limited, helping maintain the validity of the Taylor approximation in (4). By doing so, each subsystem adapts its update gate from its past history rather than the raw input token, enabling context-based selectivity: identical tokens can be filtered differently depending on their role in the sequence. Concretely, the gating vector $\Delta_{i}(k)\in\mathbb{R}^{n}$ enters the dynamics in (4) as a diagonal state feedback matrix, which modulates the contribution of the past state through $(I+A^{(i)}\mathrm{diag}(\Delta_{i}(k)))x(k-1)^{(i)},$ and the input through $\mathrm{diag}(\Delta_{i}(k))B^{(i)}u_{i}(k).$ The resulting subsystem $\Sigma^{(i)}$ is therefore context-adaptive, taking the form
$\begin{cases}x(k)^{(i)}=(I+A^{(i)}\mathrm{diag}(\Delta_{i}(k)))\,x(k-1)^{(i)}+\,\mathrm{diag}(\Delta_{i}(k))B^{(i)}\,u(k)_{i},\\ y(k)^{(i)}=C^{(i)}x(k)^{(i)}.\end{cases}$ (6)
@@
As a final step, we note that the parameterization of $(B^{(i)},C^{(i)})$ in (6) exhibits parameter redundancy: through suitable changes of basis (corresponding to simple rescalings) the same input-output behavior can be preserved while reducing the number of parameters. This fact is formalized in the following proposition, whose proof is deferred to the Appendix C.3, where a few comments are also provided on the proposition assumptions.
@@
Assume that at least one embedding vector has all the entries that are non-zero and that all the entries of the $B^{(i)}$ are non-zero. For the family of subsystems $\{\Sigma^{(i)}\}_{i=0}^{D-1}$ defined in (6), there exist changes of basis in the input and state spaces such that the structure of (6) is preserved and:
(i) all input-to-state matrices $B^{(i)}$ can be fixed to the constant vector $\mathbf{1}=[1,\dots,1]^{\top}$, with their scaling absorbed into the corresponding $C^{(i)}$, and
(ii) one embedding vector can be normalized to $\mathbf{1}$.
The transformations in Proposition 1 preserve both the input-output map and the diagonal structure of $A^{(i)}$. Consequently, the input-to-state matrices $B^{(i)}$ are redundant, as their effect can always be absorbed into $C^{(i)}$, and one embedding vector can be fixed to $\mathbf{1}$. This reduces the parameter count by $nD+D$ without loss of expressivity, yielding a leaner architecture.